John says to Brenda, "if you give me seven dollars, we will have an equal amount of money." Brenda responds, "that may be true, but if you give me seven dollars, i will have one dollar shy of having twice as much money as you." How much money did John and Brenda each have? (Show as much work as possible)
Hint: When one person gets $7 the other one loses $7.

Answer by checkley77(12844)

X=JOHN'S MONEY
Y=BRENDA'S MONEY
X+7=Y-7 OR X=Y-14
Y+7=2X-1
Y+7=2(Y-14)-1
Y+7=2Y-28-1
Y-2Y=-29-7
-Y=-36
Y=36 AMOUNT OF BRENDA'S MONEY.
X=36-14
X=22 AMOUNT OF JOHN'S MONEY.
PROOF:
36+7=2*22-1
43=44-1
43=43